### v1

# Quantum Enhancement of Thermalization

Abstract
Relaxation to equilibrium in Bose-Hubbard rings is numerically investigated in the time domain. We show that for small mean site populations in the Mott insulator regime, characterized by strong onsite interaction as compared to hopping, equipartition starting from nonequilibrium initial states is significantly faster in these bosonic quantum many-body systems than in their classical counterparts, given by discrete Gross-Pitaevskii lattices. We attribute this phenomenon of thermalization enhancement to a genuine quantum transport mechanism akin to tunneling, allowing for pathways that are more efficient than chaotic diffusion across the ergodic domain of the classical systems.
Relaxation to equilibrium has proven to be a key tool to study the thermalization and localization properties of quantum many-body systems [1, 2, 3, 4, 5, 6, 7]. State-of-the-art quantum simulation platforms using, e.g., ultracold atoms [8] or Rydberg lattices [7], allow for the preparation of a nonequilibrium initial state (either directly or via a parameter quench) as well as for the detection of the distribution of particles, spins, or excitations within the system’s configuration space after a given evolution time. An equidistribution found after a sufficiently long time is indicative of eigenstate thermalization [9] whereas its absence can often be associated with many-body localization [10] or, in specific cases, many-body scars [11].
Such relaxation studies have the additional asset that they permit comparisons with the corresponding classical dynamics of the many-body system under consideration, provided the latter exhibits a well-defined and meaningful classical counterpart. This is, in particular, the case for Bose-Hubbard (BH) systems, which describe ultracold atoms trapped in optical lattices. Their classical analog, often associated with a mean-field approximation, is given in terms of a Gross-Pitaevskii (GP) lattice. Far from being of purely academic interest, such classical studies provide complementary (and often rather intuitive) insight into the dynamical transport mechanisms that are at work in the system, and can be conducted more easily for large system sizes than the corresponding quantum calculations. Good agreement between classical and quantum dynamics is generally expected for BH systems with relatively weak interaction and/or large site populations, even though localization effects induced by quantum many-body interference can nevertheless give rise to subtle but robust obstacles towards perfect thermalization in the quantum system as opposed to its classical counterpart [12].
As we shall show here, a quantum effect of opposite nature and greater impact is to be encountered in the complementary parameter regime of BH systems, characterized by weak onsite populations and a relatively strong interaction as compared to inter-site hopping. The weakness of the latter inhibits classical transport in the framework of a GP lattice such that thermalization is attained only at very long time scales [13, 14, 15, 16]. For the corresponding quantum system, however, we show here that there are efficient and robust shortcut mechanisms akin to collective many-body tunneling, which “bridge” phase-space regions of slow transport and thus give rise to a drastic reduction of thermalization time scales as compared to what would be expected from a classical point of view.
Specifically, we numerically investigate thermalization in the evolution of Glauber coherent states in BH rings and their classical counterparts. We simulate quantum dynamics by means of a multi-configuration variational generalized coherent state approach [17], while the classical computations are performed using the truncated Wigner approximation (TWA) [18]. The efficiency of such methods allows us to reach system sizes large enough to observe relaxation towards thermal equilibrium in the time domain. Our results show that while quantum and classical systems behave similarly in the superfluid (SF) regime, characterized by weak interaction as compared to hopping, classical equipartition is severely slowed down when compared to its quantum analogue in the complementary Mott insulator (MI) regime, as showcased in Fig. 1. Since equipartition of populations is a reliable indicator for thermalization in the MI regime, this Letter describes the first numerical observation of what can be referred to as the quantum enhancement of thermalization (QET).
We start by explicitly defining the Hamiltonians describing the systems used throughout this Letter. BH rings with sites will be modelled by
where are bosonic number operators and periodic boundary conditions are assumed. The classical analog of (1) is the GP Hamiltonian
where and are the position and momentum quadratures associated to the th degree of freedom and is the dequantized number operator. The Hamiltonian in (2) can be obtained as the mean-field limit of (1) [19] or, equivalently, by a standard employment of Einstein-Brillouin-Keller theory [20, *brillouin1926remarques, *keller1958corrected].
Despite the one-to-one mapping from (1) to (2), the former describes a quantum system which, unlike the latter, will include several types of phenomena that are absent in classical physics. These genuinely quantum phenomena can be classified into two main categories depending on whether or not they can still be associated with classical trajectories in the conventional semiclassical framework. In BH systems such an association would be possible for weak localization in Fock space [12], for scars [23, 24], or for many-body revivals [25], which all arise due to constructive or destructive many-body interference caused by the superposition principle, but are perfectly amenable to quantitative reproduction via a semiclassical approximation of Feynman’s path integral [26, 27, 28, 29]. Phenomena related to barrier or dynamical tunneling [30], on the other hand, possibly enhanced by chaos and/or resonances [31, 32, 33, 34], can be considered to be intrinsically “nonclassical” as no real classical trajectory can describe the associated transition processes in phase space (at least not in an efficient manner, see [35]), and one would have to generalize the semiclassical framework to the complex domain [36, 37, 38, 39, 40] in order to properly incorporate them via complex (e.g., instanton-type) orbits.
The role that such quantum phenomena related to tunneling and interference play in many-body systems is still largely unknown, and in the following we will show that they are responsible for endowing quantum evolution with a much faster approach towards thermalization than purely classical evolution. Before moving on, however, it is important to stress that the concept of equipartition, which is employed here to serve as an experimentally accessible indicator for thermalization, does not necessarily imply the presence of the latter under all circumstances, since depending on the choice of the underlying one-body basis (which can be dictated by experimental constraints) equipartition can be observed in the presence of nearly or even fully integrable dynamics [41, 42]. For this reason, numerical studies on thermalization have been increasingly focusing on computing Lyapunov characteristic spectra (LCS) or the Kolmogorov-Sinai (KS) entropy instead of actual expectation values [43, 16], since these are invariant and coordinate-independent quantities that cannot trigger “false positives” [44]. If picking a proper observable is unavoidable, a pertinent strategy for choosing it is to focus on functions of the action variables for the nearest integrable limit, simply because right at that limit such actions will be conserved and equipartition will thus never take place (unless the system’s initial state already happens to feature identical expectation values for these actions) [45, 46]. Therefore, if equilibration of actions is still observed as the integrable limit is approached, and if it yields final equilibrium values that coincide with the ones predicted by statistical mechanics [47], then equipartition will be unavoidably linked to thermalization.
In the context of Hamiltonian (2), two integrable limits can be identified: the perfect SF, with ; and the perfect MI, where . Since the outcome of the former limit is a harmonic lattice, any classical expectation value computed from its dynamics will be identical to its quantum analogue [48]. In the latter MI regime, however, quantum expectation values will match their classical equivalents only up to the Ehrenfest time [49], rendering this regime significantly more interesting from the point of view of quantum-classical correspondence. In this limit the actions are the harmonic oscillator ones, namely . Thus, the single-site populations form the set of “good” observables in the MI regime, freezing as (but not as , in which case the actions would be given by the normal modes of the lattice). For this reason they will be used to characterize thermalization in the numerical simulations that follow.
For an optimal quantum-classical comparison of thermalization we choose initial states given by multi-mode Glauber coherent states,
which can be considered as the closest quantum analog to a classical phase-space point and which represent excellent approximations [50] for the description of Bose-Einstein condensates that are prepared on single-particle modes with the site populations . As is well-known, the dimension of the Hilbert space for this model grows quickly with both the number of lattice sites and the mean population of the system, , rendering many-site calculations with large very challenging in Fock space [51]. We thus use a variational approach, based on a multi-configuration ansatz for the solution of the time-dependent Schrödinger equation in terms of generalized coherent states, whose implementation details and convergence properties can be found in [51]. Classical TWA calculations are carried out via a Monte Carlo evaluation of the approximate expression
for the site populations. Here, is the Wigner function of the initial state (3), given by a Gaussian in phase space, and denotes the Hamiltonian flow starting from the initial phase-space point sampled according to . The sampling over several trajectories of different energies (we used trajectories in our simulations) washes out any behavior associated with atypical solutions, e.g. breathers [52, *rumpf2008transition, *mithun2018weakly], whose number is much smaller than that of the generic “thermalizing” trajectories in the regimes considered here.
Figure 1 demonstrates QET in the MI regime, for . We consider here a coherent state (3) where the populations are chosen as random numbers, drawn from a uniform distribution, such that their sum yields the total mean population in Fig. 1(a-d) and in Fig. 1(e). As can be seen in Fig. 1(b), quantum and classical time evolutions of this random Glauber state, governed by the Hamiltonians (1) and (2), respectively, eventually yield equipartition in the site populations, but on very different time scales: while the quantum calculations reach and stabilize at the thermal average of particles per site at about , their classical analogues will require times at least one order of magnitude longer in order to reach this equilibrium. To facilitate comparisons, we express the deviation from thermal equilibrium more quantitatively through the variance
of the site populations with respect to their equilibrium values, the latter being given by for a perfectly homogeneous lattice. As is seen in Figs. 1(c)–(e), QET is most clearly rendered in terms of this variance. It prevails for large system sizes (Fig. 1(d)) while it becomes less pronounced for smaller lattices with increased average site populations (Fig. 1(e) where the system is more semiclassical due to the larger ratio).
Can QET be associated with the presence of too weak chaos and thus too inefficient transport paths towards thermalization in the classical many-body system? To quantitatively address this question, we show in Fig. 2(a) the KS entropy [55] for a site system with average unit filling, , as a function of . The KS entropy is obtained from the sum of positive exponents in the LCS, which in turn is numerically computed via the prescription of [56, *geist1990comparison]. We choose for this purpose an initial phase-space point with random on-site populations, similar to the ones illustrated in Fig. 1(a), giving rise to energy densities for which the resulting classical trajectory explores a large ergodic domain ( for , for ). Unsurprisingly, the KS entropy vanishes in the two integrable SF and MI limits, while it exhibits a plateau of maximal chaos for , in agreement with the findings of [58] that were obtained through spectral studies of the BH model.
On the MI side of the plateau, the transition from well-developed chaos to near-integrability is, as is seen in Fig. 2(b)–(d), manifested in the LCS by a significant drop of the primary Lyapunov exponent as well as by a transition from a linear to an exponential decrease of the remaining positive exponents with their order [16]. Correspondingly, as shown in Fig. 2(e)–(g), chaotic transport in the classical phase space drastically slows down when going towards the MI limit, eventually giving rise to thermalization time scales that are, for , beyond numerical verification. In the corresponding quantum system, however, this transition to integrability affects thermalization to a much lesser extent. Quantum transport in the MI parameter regime can thus take place along channels that are inaccessible for the corresponding classical system, similarly as for tunneling taking place in a truly near-integrable system whose classical phase space is dominantly constituted by invariant Kolmogorov-Arnold-Moser tori. Note that these quantum channels are effective already at where the classical system is maximally chaotic, whereas they do not play any role in the SF regime, the latter featuring nearly perfect agreement between classical and quantum thermalization speeds due to exact quantum-classical correspondence [51].
Despite being of genuine quantum nature, the phenomenon of thermalization enhancement that we encounter here cannot be attributed to the occurrence of resonances in the many-body spectrum caused by accidental near-degeneracies of energy levels. To demonstrate this, we show in Fig. 3 the time evolution of the mean site populations and their corresponding variance in the presence of weak disorder. The latter is generated by randomly chosen onsite energies drawn from the uniform intervals (left column) and (right column of Fig. 3). Considering, for and , an initial coherent state with the staggered populations (which in a disorder-free context can give rise to various resonances and near-degeneracies due to its high symmetry), we clearly see that the presence of these weak random energies inhibits equilibration on a long-time scale, even when quantitatively accounting for the fact that the mean thermal site occupancies in the eigenstate thermalization hypothesis regime are no longer uniform for such a specific disorder realization [51]. This phenomenon, which is akin to many-body localization [10], can be attributed to dynamical localization [59, 60] in Fock space impeding, through destructive interference, the exploration of the system’s entire ergodic domain. Rather intriguingly, however, no significant effect of the presence of this weak disorder can be found at short and intermediate time scales, , when QET sets in. We infer from this finding that the enhancement of thermalization is a rather robust phenomenon that does not rely on fine tuning of system parameters. In spectral terms, while QET cannot be attributed to near-degeneracies of individual energy levels, we conjecture that it arises from an approximate resonance between groups of energy levels, namely those that are associated with the occupancy distribution of the initial coherent state and its counterparts obtained through permutations of the occupancies among the lattice sites. Classical transport between these groups of levels is slow for weak hopping, and quantum effects similar to tunneling allow the system to bridge across slowly diffusive phase-space regions.
Evidently, this robustness facilitates experimental verification within quantum simulation platforms based on interacting bosonic particles. In particular, the thermalization behaviour shown in Fig. 2 can be straightforwardly probed in state-of-the-art experimental setups using ultracold atoms in optical lattices [3, 4, 6] and compared to what would be expected in the corresponding classical system. Moreover, the controlled use of environmental effects [61, 62] can allow one to destroy many-body quantum coherence and effectively induce classicality (i.e., mean-field dynamics) in the experiment, thus opening possibilities for probing QET through a direct comparison of classical and quantum thermalization.
Acknowledgements.
The authors gratefully acknowledge the computing time made available to them on the high-performance computer at the NHR Center of TU Dresden. This center is jointly supported by the Federal Ministry of Education and Research and the state governments participating in the NHR (www.nhr-verein.de/unsere-partner). GML thanks Sergej Flach and Dominik Šafránek for fruitful discussions, and Technische Universität Dresden for their hospitality. GML acknowledges financial support from the Institute for Basic Science in the Republic of Korea through the project IBS-R024-D1.References
- Polkovnikov et al. [2011] A. Polkovnikov, K. Sengupta, A. Silva, and M. Vengalattore, Colloquium: Nonequilibrium dynamics of closed interacting quantum systems, Reviews of Modern Physics 83, 863 (2011).
- Trotzky et al. [2012] S. Trotzky, Y. Chen, A. Flesch, I. P. McCulloch, U. Schollwöck, J. Eisert, and I. Bloch, Probing the relaxation towards equilibrium in an isolated strongly correlated one-dimensional Bose gas, Nature Physics 8, 325 (2012).
- yoon Choi et al. [2016] J. yoon Choi, S. Hild, J. Zeiher, P. Schauß, A. Rubio-Abadal, T. Yefsah, V. Khemani, D. A. Huse, I. Bloch, and C. Gross, Exploring the many-body localization transition in two dimensions, Science 352, 1547 (2016).
- Kaufman et al. [2016] A. M. Kaufman, M. E. Tai, A. Lukin, M. Rispoli, R. Schittko, P. M. Preiss, and M. Greiner, Quantum thermalization through entanglement in an isolated many-body system, Science 353, 794 (2016).
- Reimann [2016] P. Reimann, Typical fast thermalization processes in closed many-body systems, Nature Communications 7, 10821 (2016).
- Lukin et al. [2019] A. Lukin, M. Rispoli, R. Schittko, M. E. Tai, A. M. Kaufman, S. Choi, V. Khemani, J. Léonard, and M. Greiner, Probing entanglement in a many-body-localized system, Science 364, 256 (2019).
- Bernien et al. [2017] H. Bernien, S. Schwartz, A. Keesling, H. Levine, A. Omran, H. Pichler, S. Choi, A. S. Zibrov, M. Endres, M. Greiner, V. Vuletić, and M. D. Lukin, Probing many-body dynamics on a 51-atom quantum simulator, Nature 551, 579 (2017).
- Greiner et al. [2002] M. Greiner, O. Mandel, T. Esslinger, T. W. Hänsch, and I. Bloch, Quantum phase transition from a superfluid to a Mott insulator in a gas of ultracold atoms, nature 415, 39 (2002).
- Luca D’Alessio and Rigol [2016] A. P. Luca D’Alessio, Yariv Kafri and M. Rigol, From quantum chaos and eigenstate thermalization to statistical mechanics and thermodynamics, Advances in Physics 65, 239 (2016).
- Abanin et al. [2019] D. A. Abanin, E. Altman, I. Bloch, and M. Serbyn, Colloquium: Many-body localization, thermalization, and entanglement, Rev. Mod. Phys. 91, 021001 (2019).
- Turner et al. [2018] C. J. Turner, A. A. Michailidis, D. A. Abanin, M. Serbyn, and Z. Papić, Weak ergodicity breaking from quantum many-body scars, Nature Physics 14, 745 (2018).
- Engl et al. [2014] T. Engl, J. Dujardin, A. Argüelles, P. Schlagheck, K. Richter, and J. D. Urbina, Coherent backscattering in Fock space: A signature of quantum many-body interference in interacting bosonic systems, Phys. Rev. Lett. 112, 140403 (2014).
- Livi et al. [1987] R. Livi, A. Politi, S. Ruffo, and A. Vulpiani, Liapunov exponents in high-dimensional symplectic dynamics, Journal of statistical physics 46, 147 (1987).
- Flach et al. [2009] S. Flach, D. O. Krimer, and C. Skokos, Universal spreading of wave packets in disordered nonlinear systems, Phys. Rev. Lett. 102, 024101 (2009).
- Iubini et al. [2019] S. Iubini, L. Chirondojan, G.-L. Oppo, A. Politi, and P. Politi, Dynamical freezing of relaxation to equilibrium, Physical review letters 122, 084102 (2019).
- Lando and Flach [2023] G. M. Lando and S. Flach, Thermalization slowing down in multidimensional Josephson junction networks, Physical Review E 108, L062301 (2023).
- Qiao and Grossmann [2021] Y. Qiao and F. Grossmann, Exact variational dynamics of the multimode Bose-Hubbard model based on su (m) coherent states, Physical Review A 103, 042209 (2021).
- Blakie et al. [2008] P. Blakie, A. Bradley, M. Davis, R. Ballagh, and C. Gardiner, Dynamics and statistical mechanics of ultra-cold Bose gases using c-field techniques, Advances in Physics 57, 363 (2008).
- Pitaevskii [1961] L. P. Pitaevskii, Vortex lines in an imperfect Bose gas, Sov. Phys. JETP 13, 451 (1961).
- Einstein [1917] A. Einstein, Zum Quantensatz von Sommerfeld und Epstein, Verh. d. D. Physik. Ges , 82 (1917).
- Brillouin [1926] L. Brillouin, Remarques sur la mécanique ondulatoire, J. Phys. Radium 7, 353 (1926).
- Keller [1958] J. B. Keller, Corrected Bohr-Sommerfeld quantum conditions for nonseparable systems, Annals of Physics 4, 180 (1958).
- Heller [1984] E. J. Heller, Bound-state eigenfunctions of classically chaotic hamiltonian systems: Scars of periodic orbits, Phys. Rev. Lett. 53, 1515 (1984).
- Hummel et al. [2023] Q. Hummel, K. Richter, and P. Schlagheck, Genuine many-body quantum scars along unstable modes in Bose-Hubbard systems, Phys. Rev. Lett. 130, 250402 (2023).
- Schlagheck et al. [2022] P. Schlagheck, D. Ullmo, G. M. Lando, and S. Tomsovic, Resurgent revivals in bosonic quantum gases: A striking signature of many-body quantum interferences, Phys. Rev. A 106, L051302 (2022).
- Simon and Strunz [2014] L. Simon and W. T. Strunz, Time-dependent semiclassics for ultracold bosons, Phys. Rev. A 89, 052112 (2014).
- Ray et al. [2016] S. Ray, P. Ostmann, L. Simon, F. Grossmann, and W. T. Strunz, Dynamics of interacting bosons using the Herman–Kluk semiclassical initial value representation, Journal of Physics A: Mathematical and Theoretical 49, 165303 (2016).
- Tomsovic et al. [2018] S. Tomsovic, P. Schlagheck, D. Ullmo, J.-D. Urbina, and K. Richter, Post-Ehrenfest many-body quantum interferences in ultracold atoms far out of equilibrium, Physical Review A 97, 061606 (2018).
- Lando et al. [2019] G. M. Lando, R. O. Vallejos, G.-L. Ingold, and A. M. Ozorio de Almeida, Quantum revival patterns from classical phase-space trajectories, Physical Review A 99, 042125 (2019).
- Keshavamurthy and Schlagheck [2011] S. Keshavamurthy and P. Schlagheck, eds., Dynamical Tunneling: Theory and Experiment (Taylor & Francis CRC, Boca Raton, 2011).
- Tomsovic and Ullmo [1994] S. Tomsovic and D. Ullmo, Chaos-assisted tunneling, Physical Review E 50, 145 (1994).
- Brodier et al. [2001] O. Brodier, P. Schlagheck, and D. Ullmo, Resonance-assisted tunneling in near-integrable systems, Physical Review Letters 87, 064101 (2001).
- Brodier et al. [2002] O. Brodier, P. Schlagheck, and D. Ullmo, Resonance-assisted tunneling, Annals of Physics 300, 88 (2002).
- Karmakar and Keshavamurthy [2021] S. Karmakar and S. Keshavamurthy, Arnold web and dynamical tunneling in a four-site Bose–Hubbard model, Physica D: Nonlinear Phenomena 427, 133006 (2021).
- Grossmann and Heller [1995] F. Grossmann and E. J. Heller, A semiclassical correlation function approach to barrier tunneling, Chemical Physics Letters 241, 45 – 50 (1995).
- Miller and George [1972] W. H. Miller and T. F. George, Analytic continuation of classical mechanics for classically forbidden collision processes, The Journal of Chemical Physics 56, 5668 (1972).
- Creagh [1994] S. C. Creagh, Tunnelling in multidimensional systems, Journal of Physics A: Mathematical and General 27, 4969 (1994).
- Shudo and Ikeda [1995] A. Shudo and K. S. Ikeda, Complex classical trajectories and chaotic tunneling, Phys. Rev. Lett. 74, 682 (1995).
- Takatsuka et al. [1999] K. Takatsuka, H. Ushiyama, and A. Inoue-Ushiyama, Tunneling paths in multi-dimensional semiclassical dynamics, Physics Reports 322, 347 (1999).
- Onishi et al. [2003] T. Onishi, A. Shudo, K. S. Ikeda, and K. Takahashi, Semiclassical study on tunneling processes via complex-domain chaos, Phys. Rev. E 68, 056211 (2003).
- Baldovin et al. [2021] M. Baldovin, A. Vulpiani, and G. Gradenigo, Statistical mechanics of an integrable system, Journal of Statistical Physics 183, 41 (2021).
- Baldovin et al. [2023] M. Baldovin, R. Marino, and A. Vulpiani, Ergodic observables in non-ergodic systems: The example of the harmonic chain, Physica A: Statistical Mechanics and its Applications 630, 129273 (2023).
- Malishava and Flach [2022] M. Malishava and S. Flach, Lyapunov spectrum scaling for classical many-body dynamics close to integrability, Physical Review Letters 128, 134102 (2022).
- Eichhorn et al. [2001] R. Eichhorn, S. J. Linz, and P. Hänggi, Transformation invariance of Lyapunov exponents, Chaos, Solitons & Fractals 12, 1377 (2001).
- Arnol’d [2013] V. I. Arnol’d, Mathematical methods of classical mechanics, Vol. 60 (Springer Science & Business Media, 2013).
- Ozorio de Almeida [1988] A. M. Ozorio de Almeida, Hamiltonian systems: Chaos and quantization (Cambridge University Press, 1988).
- Reimann and Evstigneev [2013] P. Reimann and M. Evstigneev, Quantum versus classical foundation of statistical mechanics under experimentally realistic conditions, Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 88, 052114 (2013).
- Feynman et al. [2010] R. P. Feynman, A. R. Hibbs, and D. F. Styer, Quantum mechanics and path integrals (Courier Corporation, 2010).
- Schubert et al. [2012] R. Schubert, R. O. Vallejos, and F. Toscano, How do wave packets spread? time evolution on Ehrenfest time scales, Journal of Physics A: Mathematical and Theoretical 45, 215307 (2012).
- Lieb et al. [2007] E. H. Lieb, R. Seiringer, and J. Yngvason, Bose-Einstein condensation and spontaneous symmetry breaking, Reports on Mathematical Physics 59, 389 (2007).
- [51] See Supplemental Material, which includes references [63, 64, 17, 65], at [url will be inserted by publisher] for a comprehensive account of numerical and analytical details including convergence of quantum simulations, the calculation of thermal averages in the disordered rings, and comparisons performed in the superfluid regime, .
- Flach and Willis [1998] S. Flach and C. R. Willis, Discrete breathers, Physics reports 295, 181 (1998).
- Rumpf [2008] B. Rumpf, Transition behavior of the discrete nonlinear Schrödinger equation, Physical Review E 77, 036606 (2008).
- Mithun et al. [2018] T. Mithun, Y. Kati, C. Danieli, and S. Flach, Weakly nonergodic dynamics in the Gross-Pitaevskii lattice, Physical review letters 120, 184101 (2018).
- Cassidy et al. [2009] A. C. Cassidy, D. Mason, V. Dunjko, and M. Olshanii, Threshold for Chaos and Thermalization in the One-Dimensional Mean-Field Bose-Hubbard Model, Phys. Rev. Lett. 102, 025302 (2009).
- Benettin et al. [1980] G. Benettin, L. Galgani, A. Giorgilli, and J.-M. Strelcyn, Lyapunov characteristic exponents for smooth dynamical systems and for Hamiltonian systems: a method for computing all of them. Parts 1 and 2, Meccanica 15, 9 (1980).
- Geist et al. [1990] K. Geist, U. Parlitz, and W. Lauterborn, Comparison of different methods for computing Lyapunov exponents, Prog. Theor. Phys. 83, 875 (1990).
- Pausch et al. [2022] L. Pausch, A. Buchleitner, E. G. Carnio, and A. Rodríguez, Optimal route to quantum chaos in the Bose–Hubbard model, Journal of Physics A: Mathematical and Theoretical 55, 324002 (2022).
- Fishman et al. [1982] S. Fishman, D. R. Grempel, and R. E. Prange, Chaos, quantum recurrences, and anderson localization, Phys. Rev. Lett. 49, 509 (1982).
- Ishikawa et al. [2007] A. Ishikawa, A. Tanaka, and A. Shudo, Quantum suppression of chaotic tunnelling, Journal of Physics A: Mathematical and Theoretical 40, F397 (2007).
- Witthaut et al. [2008] D. Witthaut, F. Trimborn, and S. Wimberger, Dissipation induced coherence of a two-mode Bose-Einstein condensate, Phys. Rev. Lett. 101, 200402 (2008).
- Barontini et al. [2013] G. Barontini, R. Labouvie, F. Stubenrauch, A. Vogler, V. Guarrera, and H. Ott, Controlling the dynamics of an open many-body quantum system with localized dissipation, Phys. Rev. Lett. 110, 035302 (2013).
- Buonsante and Penna [2008] P. Buonsante and V. Penna, Some remarks on the coherent-state variational approach to nonlinear boson models, Journal of Physics A: Mathematical and Theoretical 41, 175301 (2008).
- Trimborn et al. [2008] F. Trimborn, D. Witthaut, and H. J. Korsch, Exact number-conserving phase-space dynamics of the -site Bose-Hubbard model, Phys. Rev. A 77, 043631 (2008).
- Qiao and Grossmann [2023] Y. Qiao and F. Grossmann, Revealing quantum effects in bosonic Josephson junctions: a multi-configuration atomic coherent state approach, Front. Phys. 11, 1221614 (2023).