### v2

# Quantum Enhancement of Thermalization

Abstract
Equilibrium properties of many-body systems with a large number of degrees of freedom are generally expected to be described by statistical mechanics. Such expectation is closely tied to the observation of thermalization, as manifested through equipartition in time-dependent observables, which takes place both in quantum and classical systems but may look very different in comparison. By studying the dynamics of individual lattice site populations in ultracold bosonic gases, it is shown that the process of relaxation towards equilibrium in a quantum system can be orders of magnitude faster than in its classical counterpart. Classical chaos quantifiers reveal that this is due to a wave packet in a quantum system being able to escape regions of inefficient classical transport by a mechanism akin to tunneling. Since the presented phenomenon takes place in a broad parameter range and persists in weakly disordered systems, we expect that it occurs in a variety of many-body systems and is amenable to direct experimental verification in state-of-the-art quantum simulation platforms.
Relaxation to equilibrium has proven to be a key tool for studying the thermalization and localization properties of quantum many-body systems [1, 2, 3, 4, 5, 6, 7]. State-of-the-art quantum simulation platforms using, e.g., ultracold atoms [8] or Rydberg lattices [7], allow for the preparation of a non-equilibrium initial state (either directly or via a parameter quench) as well as for the detection of the distribution of particles, spins, or excitations within the system’s configuration space after a given evolution time. An equidistribution found after a sufficiently long time is indicative of eigenstate thermalization [9] whereas its absence can often be associated with many-body localization [10] or, in specific cases, many-body scars [11].
Such relaxation studies have the additional asset that they permit comparisons with the corresponding classical dynamics of the many-body system under consideration, provided the latter exhibits a well-defined and meaningful classical counterpart. This is, in particular, the case for Bose-Hubbard (BH) systems, which describe ultracold atoms trapped in optical lattices. Their classical analogue, often associated with a mean-field approximation, are Gross-Pitaevskii (GP) lattices. Studying such classical analogues provides complementary (and often intuitive) insight into the dynamical transport mechanisms that are at work during evolution, and can be conducted more easily for large system sizes than the corresponding quantum calculations. Good agreement between classical and quantum dynamics is generally expected for BH systems with relatively weak interaction and/or large site populations, even though localization effects induced by quantum many-body interference can nevertheless give rise to subtle but robust obstacles toward perfect thermalization in the quantum system as opposed to its classical counterpart [12, 13].
As we shall show here, a quantum effect of opposite nature and greater impact is to be encountered in the complementary parameter regime of BH systems, characterized by small onsite populations and a relatively strong interaction as compared to inter-site hopping. The weakness of the latter inhibits classical transport in the framework of a GP lattice such that thermalization is attained only at very long time scales [14, 15, 16, 17]. For the corresponding quantum system, however, we show here that there are efficient and robust shortcut mechanisms akin to collective many-body tunneling, which bridge phase-space regions of slow transport and thus give rise to a drastic reduction of thermalization time scales as compared to what would be expected from a classical point of view. Since such a slowdown in classical diffusion is a property shared by several many-body systems in the weakly coupled regime, we expect this phenomenon to be more general than for BH systems alone.
Specifically, we numerically investigate thermalization in the evolution of Glauber coherent states in BH rings and their classical counterparts. We simulate quantum dynamics by means of a multi-configuration variational generalized coherent state approach [18, 19], while the classical computations are performed using the truncated Wigner approximation [20, 21]. The efficiency of such methods allows us to reach system sizes large enough to observe relaxation toward thermal equilibrium in the time domain. Our results show that while quantum and classical systems behave similarly in the weakly interacting regime, where the onsite interactions are small when compared to hopping, classical equipartition is severely slowed down when compared to its quantum analogue in the complementary strongly interacting case, as showcased in Fig. 1. Since equipartition of populations is a reliable indicator of thermalization in this latter regime, this Letter describes the first numerical observation of what can be referred to as the quantum enhancement of thermalization (QET).
We start by explicitly writing the quantum and classical Hamiltonians describing the systems used throughout this Letter. BH rings with sites will be modeled by
where periodic boundary conditions are assumed and , with () representing bosonic creation (annihilation) operators. The first term above represents onsite interactions, with strength , while the second term models hopping between sites, with strength . The classical analog of (1) is the GP Hamiltonian
where and are the position and momentum quadratures associated to the th degree of freedom and is its dequantized number operator. The Hamiltonian in (2) can be obtained from (1) as a mean-field approximation [22]. The many-body wave function is then approximated by a classical field known as the order parameter, , whose dynamics are described by a discrete nonlinear Schrödinger equation. This equation is identical to Hamilton’s equations obtained from (2).
Despite the one-to-one mapping from (1) to (2), the former describes a quantum system which, unlike the latter, presents several types of phenomena that are absent in classical physics. These genuinely quantum phenomena can be classified into two main categories depending on whether or not they can still be associated with classical trajectories in the conventional semiclassical framework. In BH systems such an association would be possible for weak localization in Fock space [13], for scars [23, 24], or for many-body revivals [25], which all arise due to constructive or destructive many-body interference caused by the superposition principle, but are perfectly amenable to quantitative semiclassical reproduction [26, 27, 28, 29]. Phenomena related to barrier or dynamical tunneling [30], on the other hand, possibly enhanced by chaos and/or resonances [31, 32, 33, 34, 35, 36], can be considered to be intrinsically “nonclassical” as no real classical trajectory can describe the associated transition processes in phase space [37], and one would have to generalize the semiclassical framework to the complex domain in order to properly incorporate them via complex (e.g., instanton-type) orbits [38, 39, 40, 41, 42].
The role that such phenomena play in the statistical mechanics of quantum many-body systems is still largely unknown, and in the following we will show that they are responsible for endowing quantum evolution with a much faster approach towards thermalization than purely classical evolution. Before moving on, however, it is important to stress that the concept of equipartition, which is employed here to serve as an experimentally accessible indicator for thermalization, does not necessarily imply the presence of the latter under all circumstances, since depending on the choice of the underlying one-body basis (which can be dictated by experimental constraints) equipartition can be observed in the presence of nearly or even fully integrable dynamics [43, 44]. For this reason, numerical studies on thermalization have been increasingly focusing on computing Lyapunov spectra or the Kolmogorov-Sinai entropy instead of expectation values [45, 17], since these are invariant and coordinate-independent quantities that cannot trigger “false positives” [46]. If picking an observable is unavoidable, a pertinent strategy for choosing it is to focus on functions of the action variables for the nearest integrable limit, simply because right at that limit such actions will be conserved and equipartition cannot take place [47, 48]. Thus, if equipartition of such functions persists while approaching integrability and matches statistical mechanics predictions [49], it will be necessarily linked to thermalization.
In the context of Hamiltonian (2), two integrable limits can be identified: the harmonic limit, with ; and the limit of decoupled sites, where . Due to the near exact quantum-classical correspondence, classical expectation values lie close to their quantum analogues in the neighborhood of the harmonic limit [50]. For nearly decoupled sites, however, quantum expectation values will only match their classical counterparts up to Ehrenfest time [51]. Since the actions in this case are the harmonic oscillator ones, namely , single-site populations form the set of “good” observables in the strongly interacting regime, freezing as (but not as , in which case the actions are the normal modes of the lattice). For this reason they will be used to characterize thermalization in the numerical simulations that follow.
For an optimal quantum-classical comparison of thermalization we choose initial states given by multi-mode Glauber coherent states,
where is the vacuum state. The coherent state above is the closest quantum analog to a classical phase-space point and is an excellent approximation of Bose-Einstein condensates prepared on single-particle modes with site populations [52], which are often used as initial states in quench experiments [53]. As is well known, the dimension of the Hilbert space for the BH model grows quickly with both the number of lattice sites and the mean population of the system, , rendering many-site calculations with large very challenging in Fock space. We thus use a variational approach, based on a multi-layer extension of the multi-configuration ansatz in terms of a linear combination of time-dependent generalized coherent states with fixed particle number, introduced in [18]. Here, a summation over a range of particle numbers of the generalized coherent states, defined as , is performed in addition to the summation over the ansatz multiplicity, as explained in detail in [54].
Classical results are obtained via a Monte Carlo evaluation of the expression
which is the truncated Wigner approximation for the site populations [21]. Here, is the Wigner function of the initial state (3), given by a Gaussian in -dimensional phase space, and denotes the Hamiltonian flow of (2) starting from the initial phase-space point , which is sampled according to . Sampling over several mean-field trajectories of different energies (we used trajectories in our simulations) washes out any behavior associated with atypical solutions, e.g. breathers [55, *rumpf2008transition, *mithun2018weakly], whose number is much smaller than that of the generic “thermalizing” trajectories in the regimes considered here.
Figure 1 demonstrates QET in the regime of strong interactions, with . We consider here a coherent state (3) where the populations are chosen as random numbers drawn from a uniform distribution, such that their sum yields the total mean population in Fig. 1(a-d) and in Fig. 1(e). As can be seen in Fig. 1(b), quantum and classical time evolutions eventually yield equipartition in the site populations, but on very different time scales: while the quantum calculations reach and stabilize at the thermal average of particles per site at about , their classical analogues will require times at least one order of magnitude longer in order to reach equilibrium. To facilitate comparisons, we express the deviation from thermal equilibrium more quantitatively through the variance
of the site populations with respect to their equilibrium values, namely for a perfectly homogeneous lattice. As is seen in Figs. 1(c)–(e), QET is most clearly rendered in terms of this variance. It prevails for large system sizes (Fig. 1(d)) while it becomes less pronounced for smaller lattices with increased average site populations (Fig. 1(e) where the system is more semiclassical due to the larger ratio).
How is QET affected by the degree of chaos, which impacts the efficiency of chaotic transport paths towards thermalization in the classical many-body system? To quantitatively address this question, we show in Fig. 2(a) the Kolmogorov-Sinai entropy [58], , for a site system with average unit filling, , as a function of . Such entropy is obtained from the sum of positive exponents in the Lyapunov spectrum, computed via the prescription of [59, *geist1990comparison]. We choose for this purpose an initial phase-space point with random on-site populations 111We have verified that the Lyapunov spectra and their associated Kolmogorov-Sinai entropies are quantitatively similar for several phase-space points sampled according to the Wigner function of the initial state., giving rise to energy densities for which the resulting classical trajectory explores a large domain in phase space ( for , for ). Unsurprisingly, vanishes in the harmonic and decoupled integrable limits, while it exhibits a plateau of maximal chaos for , in agreement with a previous study on spectral properties of the BH model [62]. QET shown in Fig. 1 is thus occurring at highly developed chaos, far away from any near-integrable limit.
On the strongly interacting side of the plateau, the transition from well-developed chaos to near-integrability is, as is seen in Fig. 2(b)–(d), manifested in the Lyapunov spectrum by a significant drop in its maximal exponent as well as by a transition from a linear to an exponential decrease of the remaining positive exponents with their index, [17]. The development of near-zero exponents indicates that reducing the coupling strength gives birth to multiple near-conserved quantities. Correspondingly, as shown in Fig. 2(e)–(g), chaotic transport in the classical phase space drastically slows down in this regime, eventually giving rise to thermalization time scales that are, for , beyond numerical verification. In the corresponding quantum system, however, this proximity to integrability affects thermalization to a much lesser extent. Quantum transport in the regime of strong interactions can thus take place along channels that are inaccessible for the corresponding classical system, similarly to tunneling taking place in a truly near-integrable system whose classical phase space is dominantly constituted by invariant Kolmogorov-Arnold-Moser tori. Note that these quantum channels are already effective at where the classical system is maximally chaotic, whereas they do not play any role in the weak interaction regime, which features nearly perfect agreement between the classical and quantum thermalization speeds [54]. We note that such regimes of approximate harmonicity or nearly decoupled sites exist in a plethora of models that display the same characteristics in their Lyapunov spectra as seen in Fig. 2, such as Josephson junction arrays [17], in which QET should also take place.
Despite being of genuine quantum nature, the phenomenon of thermalization enhancement that we encounter here cannot be attributed to the occurrence of resonances in the many-body spectrum caused by accidental near-degeneracies of energy levels. To demonstrate this, we show in Fig. 3 the time evolution of the mean site populations and their corresponding variance in the presence of weak disorder. The latter is generated by randomly chosen onsite energies drawn from the uniform intervals (left column) and (right column of Fig. 3). Considering, for and , an initial coherent state with the staggered populations (which in a disorder-free context can give rise to various resonances and near-degeneracies due to its high symmetry), we clearly see that the presence of these weak random energies inhibits equilibration on a long-time scale, even when quantitatively accounting for the fact that the mean thermal site occupancies are no longer uniform for such a specific disorder realization [54]. This slowing down of thermalization is attributed to the disordered system possessing a large number of localized (nonthermal) eigenstates that violate the eigenstate thermalization hypothesis [63, 64]. Dynamical localization in Fock space then hinders the exploration of the system’s entire chaotic domain through destructive interference [65, 66].
Rather intriguingly, however, no significant effect of the presence of weak disorder can be found at short and intermediate time scales, , when QET sets in. We infer from this finding that the enhancement of thermalization is a rather robust phenomenon that does not rely on fine tuning of system parameters. In spectral terms, while QET cannot be attributed to degeneracies of individual energy levels, we conjecture that it arises from an approximate resonance between groups of energy levels, namely those that are associated with the occupancy distribution of the initial coherent state and its counterparts obtained through permutations of the occupancies among the lattice sites. Robustness with respect to system parameter variations is then granted by the multitude of individual site distribution states belonging to these groups, in combination with the fact that the effective coupling between these states can be relatively large, owing to the absence of true tunneling barriers [34, 35] (as is indeed supported by numerical findings [54]). Hence, even if in the presence of disorder only a few of these states effectively feature an approximate near-degeneracy with components of the initial coherent state, these remaining “escape channels” will be sufficient for the quantum system to bridge across slowly diffusive phase-space regions 222Note that in the strong hopping regime this mechanism is expected to become ineffective as compared to transport via classical channels, in close analogy to chaos-assisted tunneling suppressing ordinary tunneling channels between symmetric wells [31]..
We expect QET to occur in a variety of physically relevant quantum many-body systems in weakly coupled regimes, ranging from complex molecules and spin chains [68, 69] to quantum computing and simulation [7, 70]. In particular, the thermalization behavior shown in Fig. 2 can be probed in state-of-the-art experimental setups using ultracold bosonic atoms in optical lattices [3, 4, 6]. Here, the controlled use of environmental effects [71, 72] can allow one to destroy many-body quantum coherence and effectively induce classicality (i.e., mean-field dynamics) in the experiment, thus opening possibilities for probing QET through a direct comparison of classical and quantum equipartition speeds.
Acknowledgements.
The authors gratefully acknowledge the computing time made available to them on the high-performance computer at the NHR Center of TU Dresden. This center is jointly supported by the Federal Ministry of Education and Research and the state governments participating in the NHR (www.nhr-verein.de/unsere-partner). GML thanks Sergej Flach and Dominik Šafránek for fruitful discussions, and Technische Universität Dresden for their hospitality. GML acknowledges financial support from the Institute for Basic Science in the Republic of Korea through the project IBS-R024-D1.References
- Polkovnikov et al. [2011] A. Polkovnikov, K. Sengupta, A. Silva, and M. Vengalattore, Colloquium: Nonequilibrium dynamics of closed interacting quantum systems, Reviews of Modern Physics 83, 863 (2011).
- Trotzky et al. [2012] S. Trotzky, Y. Chen, A. Flesch, I. P. McCulloch, U. Schollwöck, J. Eisert, and I. Bloch, Probing the relaxation towards equilibrium in an isolated strongly correlated one-dimensional Bose gas, Nature Physics 8, 325 (2012).
- yoon Choi et al. [2016] J. yoon Choi, S. Hild, J. Zeiher, P. Schauß, A. Rubio-Abadal, T. Yefsah, V. Khemani, D. A. Huse, I. Bloch, and C. Gross, Exploring the many-body localization transition in two dimensions, Science 352, 1547 (2016).
- Kaufman et al. [2016] A. M. Kaufman, M. E. Tai, A. Lukin, M. Rispoli, R. Schittko, P. M. Preiss, and M. Greiner, Quantum thermalization through entanglement in an isolated many-body system, Science 353, 794 (2016).
- Reimann [2016] P. Reimann, Typical fast thermalization processes in closed many-body systems, Nature Communications 7, 10821 (2016).
- Lukin et al. [2019] A. Lukin, M. Rispoli, R. Schittko, M. E. Tai, A. M. Kaufman, S. Choi, V. Khemani, J. Léonard, and M. Greiner, Probing entanglement in a many-body-localized system, Science 364, 256 (2019).
- Bernien et al. [2017] H. Bernien, S. Schwartz, A. Keesling, H. Levine, A. Omran, H. Pichler, S. Choi, A. S. Zibrov, M. Endres, M. Greiner, V. Vuletić, and M. D. Lukin, Probing many-body dynamics on a 51-atom quantum simulator, Nature 551, 579 (2017).
- Greiner et al. [2002a] M. Greiner, O. Mandel, T. Esslinger, T. W. Hänsch, and I. Bloch, Quantum phase transition from a superfluid to a Mott insulator in a gas of ultracold atoms, nature 415, 39 (2002a).
- Luca D’Alessio and Rigol [2016] A. P. Luca D’Alessio, Yariv Kafri and M. Rigol, From quantum chaos and eigenstate thermalization to statistical mechanics and thermodynamics, Advances in Physics 65, 239 (2016).
- Abanin et al. [2019] D. A. Abanin, E. Altman, I. Bloch, and M. Serbyn, Colloquium: Many-body localization, thermalization, and entanglement, Rev. Mod. Phys. 91, 021001 (2019).
- Turner et al. [2018] C. J. Turner, A. A. Michailidis, D. A. Abanin, M. Serbyn, and Z. Papić, Weak ergodicity breaking from quantum many-body scars, Nature Physics 14, 745 (2018).
- Richaud and Penna [2018] A. Richaud and V. Penna, Phase separation can be stronger than chaos, New Journal of Physics 20, 105008 (2018).
- Engl et al. [2014] T. Engl, J. Dujardin, A. Argüelles, P. Schlagheck, K. Richter, and J. D. Urbina, Coherent backscattering in Fock space: A signature of quantum many-body interference in interacting bosonic systems, Phys. Rev. Lett. 112, 140403 (2014).
- Livi et al. [1987] R. Livi, A. Politi, S. Ruffo, and A. Vulpiani, Liapunov exponents in high-dimensional symplectic dynamics, Journal of statistical physics 46, 147 (1987).
- Flach et al. [2009] S. Flach, D. O. Krimer, and C. Skokos, Universal spreading of wave packets in disordered nonlinear systems, Phys. Rev. Lett. 102, 024101 (2009).
- Iubini et al. [2019] S. Iubini, L. Chirondojan, G.-L. Oppo, A. Politi, and P. Politi, Dynamical freezing of relaxation to equilibrium, Phys. Rev. Lett. 122, 084102 (2019).
- Lando and Flach [2023] G. M. Lando and S. Flach, Thermalization slowing down in multidimensional Josephson junction networks, Physical Review E 108, L062301 (2023).
- Qiao and Grossmann [2021] Y. Qiao and F. Grossmann, Exact variational dynamics of the multimode Bose-Hubbard model based on SU() coherent states, Physical Review A 103, 042209 (2021).
- Qiao [2024] Y. Qiao, Dynamical multi-configuration generalized coherent states approach to many-body bosonic quantum systems, PhD thesis, Technische Universität Dresden, Dresden, Germany (2024), available at https://nbn-resolving.org/urn:nbn:de:bsz:14-qucosa2-921049.
- Blakie et al. [2008] P. Blakie, A. Bradley, M. Davis, R. Ballagh, and C. Gardiner, Dynamics and statistical mechanics of ultra-cold Bose gases using c-field techniques, Advances in Physics 57, 363 (2008).
- Steel et al. [1998] M. J. Steel, M. Olsen, L. Plimak, P. Drummond, S. Tan, M. Collett, D. Walls, and R. Graham, Dynamical quantum noise in trapped bose-einstein condensates, Physical Review A 58, 4824 (1998).
- Richter et al. [2022] K. Richter, J. D. Urbina, and S. Tomsovic, Semiclassical roots of universality in many-body quantum chaos, Journal of Physics A: Mathematical and Theoretical 55, 453001 (2022).
- Heller [1984] E. J. Heller, Bound-state eigenfunctions of classically chaotic hamiltonian systems: Scars of periodic orbits, Phys. Rev. Lett. 53, 1515 (1984).
- Hummel et al. [2023] Q. Hummel, K. Richter, and P. Schlagheck, Genuine many-body quantum scars along unstable modes in Bose-Hubbard systems, Phys. Rev. Lett. 130, 250402 (2023).
- Schlagheck et al. [2022] P. Schlagheck, D. Ullmo, G. M. Lando, and S. Tomsovic, Resurgent revivals in bosonic quantum gases: A striking signature of many-body quantum interferences, Phys. Rev. A 106, L051302 (2022).
- Simon and Strunz [2014] L. Simon and W. T. Strunz, Time-dependent semiclassics for ultracold bosons, Phys. Rev. A 89, 052112 (2014).
- Ray et al. [2016] S. Ray, P. Ostmann, L. Simon, F. Grossmann, and W. T. Strunz, Dynamics of interacting bosons using the Herman–Kluk semiclassical initial value representation, Journal of Physics A: Mathematical and Theoretical 49, 165303 (2016).
- Tomsovic et al. [2018] S. Tomsovic, P. Schlagheck, D. Ullmo, J.-D. Urbina, and K. Richter, Post-Ehrenfest many-body quantum interferences in ultracold atoms far out of equilibrium, Physical Review A 97, 061606 (2018).
- Lando et al. [2019] G. M. Lando, R. O. Vallejos, G.-L. Ingold, and A. M. Ozorio de Almeida, Quantum revival patterns from classical phase-space trajectories, Physical Review A 99, 042125 (2019).
- Keshavamurthy and Schlagheck [2011] S. Keshavamurthy and P. Schlagheck, eds., Dynamical Tunneling: Theory and Experiment (Taylor & Francis CRC, Boca Raton, 2011).
- Tomsovic and Ullmo [1994] S. Tomsovic and D. Ullmo, Chaos-assisted tunneling, Physical Review E 50, 145 (1994).
- Brodier et al. [2001] O. Brodier, P. Schlagheck, and D. Ullmo, Resonance-assisted tunneling in near-integrable systems, Physical Review Letters 87, 064101 (2001).
- Brodier et al. [2002] O. Brodier, P. Schlagheck, and D. Ullmo, Resonance-assisted tunneling, Annals of Physics 300, 88 (2002).
- Löck et al. [2010] S. Löck, A. Bäcker, R. Ketzmerick, and P. Schlagheck, Regular-to-chaotic tunneling rates: From the quantum to the semiclassical regime, Phys. Rev. Lett. 104, 114101 (2010).
- Schlagheck et al. [2011] P. Schlagheck, A. Mouchet, and D. Ullmo, Resonance-assisted tunneling in mixed regular–chaotic systems, in Dynamical Tunneling: Theory and Experiment, edited by S. Keshavamurthy and P. Schlagheck (Taylor & Francis CRC, 2011).
- Karmakar and Keshavamurthy [2021] S. Karmakar and S. Keshavamurthy, Arnold web and dynamical tunneling in a four-site Bose–Hubbard model, Physica D: Nonlinear Phenomena 427, 133006 (2021).
- Grossmann and Heller [1995] F. Grossmann and E. J. Heller, A semiclassical correlation function approach to barrier tunneling, Chemical Physics Letters 241, 45 – 50 (1995).
- Miller and George [1972] W. H. Miller and T. F. George, Analytic continuation of classical mechanics for classically forbidden collision processes, The Journal of Chemical Physics 56, 5668 (1972).
- Creagh [1994] S. C. Creagh, Tunnelling in multidimensional systems, Journal of Physics A: Mathematical and General 27, 4969 (1994).
- Shudo and Ikeda [1995] A. Shudo and K. S. Ikeda, Complex classical trajectories and chaotic tunneling, Phys. Rev. Lett. 74, 682 (1995).
- Takatsuka et al. [1999] K. Takatsuka, H. Ushiyama, and A. Inoue-Ushiyama, Tunneling paths in multi-dimensional semiclassical dynamics, Physics Reports 322, 347 (1999).
- Onishi et al. [2003] T. Onishi, A. Shudo, K. S. Ikeda, and K. Takahashi, Semiclassical study on tunneling processes via complex-domain chaos, Phys. Rev. E 68, 056211 (2003).
- Baldovin et al. [2021] M. Baldovin, A. Vulpiani, and G. Gradenigo, Statistical mechanics of an integrable system, Journal of Statistical Physics 183, 41 (2021).
- Baldovin et al. [2023] M. Baldovin, R. Marino, and A. Vulpiani, Ergodic observables in non-ergodic systems: The example of the harmonic chain, Physica A: Statistical Mechanics and its Applications 630, 129273 (2023).
- Malishava and Flach [2022] M. Malishava and S. Flach, Lyapunov spectrum scaling for classical many-body dynamics close to integrability, Physical Review Letters 128, 134102 (2022).
- Eichhorn et al. [2001] R. Eichhorn, S. J. Linz, and P. Hänggi, Transformation invariance of Lyapunov exponents, Chaos, Solitons & Fractals 12, 1377 (2001).
- Arnol’d [2013] V. I. Arnol’d, Mathematical methods of classical mechanics, Vol. 60 (Springer Science & Business Media, 2013).
- Ozorio de Almeida [1988] A. M. Ozorio de Almeida, Hamiltonian systems: Chaos and quantization (Cambridge University Press, 1988).
- Reimann and Evstigneev [2013] P. Reimann and M. Evstigneev, Quantum versus classical foundation of statistical mechanics under experimentally realistic conditions, Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 88, 052114 (2013).
- Feynman et al. [2010] R. P. Feynman, A. R. Hibbs, and D. F. Styer, Quantum mechanics and path integrals (Courier Corporation, 2010).
- Schubert et al. [2012] R. Schubert, R. O. Vallejos, and F. Toscano, How do wave packets spread? time evolution on Ehrenfest time scales, Journal of Physics A: Mathematical and Theoretical 45, 215307 (2012).
- Lieb et al. [2007] E. H. Lieb, R. Seiringer, and J. Yngvason, Bose-Einstein condensation and spontaneous symmetry breaking, Reports on Mathematical Physics 59, 389 (2007).
- Greiner et al. [2002b] M. Greiner, O. Mandel, T. W. Hänsch, and I. Bloch, Collapse and revival of the matter wave field of a Bose–Einstein condensate, Nature 419, 51 (2002b).
- [54] See Supplemental Material, which includes references [73, 74, 18, 75, 76], at [url will be inserted by publisher] for a comprehensive account of numerical and analytical details including convergence of quantum simulations, the calculation of thermal averages in the disordered rings, comparisons performed in the regime of weak interactions, and a study of transition probabilities between near resonant states that sheds light on the mechanism of QET.
- Flach and Willis [1998] S. Flach and C. R. Willis, Discrete breathers, Physics reports 295, 181 (1998).
- Rumpf [2008] B. Rumpf, Transition behavior of the discrete nonlinear Schrödinger equation, Physical Review E 77, 036606 (2008).
- Mithun et al. [2018] T. Mithun, Y. Kati, C. Danieli, and S. Flach, Weakly nonergodic dynamics in the Gross-Pitaevskii lattice, Physical review letters 120, 184101 (2018).
- Cassidy et al. [2009] A. C. Cassidy, D. Mason, V. Dunjko, and M. Olshanii, Threshold for Chaos and Thermalization in the One-Dimensional Mean-Field Bose-Hubbard Model, Phys. Rev. Lett. 102, 025302 (2009).
- Benettin et al. [1980] G. Benettin, L. Galgani, A. Giorgilli, and J.-M. Strelcyn, Lyapunov characteristic exponents for smooth dynamical systems and for Hamiltonian systems: a method for computing all of them. Parts 1 and 2, Meccanica 15, 9 (1980).
- Geist et al. [1990] K. Geist, U. Parlitz, and W. Lauterborn, Comparison of different methods for computing Lyapunov exponents, Prog. Theor. Phys. 83, 875 (1990).
- Note [1] We have verified that the Lyapunov spectra and their associated Kolmogorov-Sinai entropies are quantitatively similar for several phase-space points sampled according to the Wigner function of the initial state.
- Pausch et al. [2022] L. Pausch, A. Buchleitner, E. G. Carnio, and A. Rodríguez, Optimal route to quantum chaos in the Bose–Hubbard model, Journal of Physics A: Mathematical and Theoretical 55, 324002 (2022).
- Deutsch [1991] J. M. Deutsch, Quantum statistical mechanics in a closed system, Physical review a 43, 2046 (1991).
- Srednicki [1994] M. Srednicki, Chaos and quantum thermalization, Physical Review E 50, 888 (1994).
- Fishman et al. [1982] S. Fishman, D. R. Grempel, and R. E. Prange, Chaos, quantum recurrences, and anderson localization, Phys. Rev. Lett. 49, 509 (1982).
- Ishikawa et al. [2007] A. Ishikawa, A. Tanaka, and A. Shudo, Quantum suppression of chaotic tunnelling, Journal of Physics A: Mathematical and Theoretical 40, F397 (2007).
- Note [2] Note that in the strong hopping regime this mechanism is expected to become ineffective as compared to transport via classical channels, in close analogy to chaos-assisted tunneling suppressing ordinary tunneling channels between symmetric wells [31].
- Bardet et al. [2023] I. Bardet, Á. Capel, L. Gao, A. Lucia, D. Pérez-García, and C. Rouzé, Rapid thermalization of spin chain commuting hamiltonians, Physical Review Letters 130, 060401 (2023).
- Kumar and Choudhury [2024] S. Kumar and S. Choudhury, Prethermalization in aperiodically driven classical spin systems, arXiv preprint arXiv:2404.10224 (2024).
- Roushan et al. [2017] P. Roushan, C. Neill, J. Tangpanitanon, V. M. Bastidas, A. Megrant, R. Barends, Y. Chen, Z. Chen, B. Chiaro, A. Dunsworth, A. Fowler, B. Foxen, M. Giustina, E. Jeffrey, J. Kelly, E. Lucero, J. Mutus, M. Neeley, C. Quintana, D. Sank, A. Vainsencher, J. Wenner, T. White, H. Neven, D. G. Angelakis, and J. Martinis, Spectroscopic signatures of localization with interacting photons in superconducting qubits, Science 358, 1175 (2017).
- Witthaut et al. [2008] D. Witthaut, F. Trimborn, and S. Wimberger, Dissipation induced coherence of a two-mode Bose-Einstein condensate, Phys. Rev. Lett. 101, 200402 (2008).
- Barontini et al. [2013] G. Barontini, R. Labouvie, F. Stubenrauch, A. Vogler, V. Guarrera, and H. Ott, Controlling the dynamics of an open many-body quantum system with localized dissipation, Phys. Rev. Lett. 110, 035302 (2013).
- Buonsante and Penna [2008] P. Buonsante and V. Penna, Some remarks on the coherent-state variational approach to nonlinear boson models, Journal of Physics A: Mathematical and Theoretical 41, 175301 (2008).
- Trimborn et al. [2008] F. Trimborn, D. Witthaut, and H. J. Korsch, Exact number-conserving phase-space dynamics of the -site Bose-Hubbard model, Phys. Rev. A 77, 043631 (2008).
- Dell’Anna [2022] L. Dell’Anna, Entanglement properties and ground-state statistics of free bosons, Phys. Rev. A 105, 032412 (2022).
- Qiao and Grossmann [2023] Y. Qiao and F. Grossmann, Revealing quantum effects in bosonic Josephson junctions: a multi-configuration atomic coherent state approach, Front. Phys. 11, 1221614 (2023).